Indian Airlines have a capacity to carry a maximum of 10.000 passengers per month from

Calcutta to Guhawati at a fare of Rs. 500. Variable costs are Rs. 100 per passenger and fixed

costs are Rs. 30,000 per month. How many passengers should be carried per month to break

even


1
Expert's answer
2022-01-26T10:00:20-0500

Solution:

Break-even units = Fixed Costs ÷ Contribution Margin

Contribution Margin = Revenue per unit – Variable Costs per unit = 500 – 100 = 400

Fixed Costs = 30,000


Break-even units = 30,000 ÷ 400 = 75 passengers


To break even, 75 passengers per month should be carried.